OpenStudy (anonymous):

can someone please show me how to do this? solve the equation: 2cosx+1=0

OpenStudy (harsimran_hs4):

cos x = -1/2 so can you tell me the values of x satisfying the above eq in the range[0, 2pi)?

OpenStudy (saifoo.khan):

\[2 \cos x +1 = 0\]\[2 \cos x =-1 \]Divide 2 on both sides, \[\cos x = -\frac12\]Then follow above^

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

So why are you reposting this?

OpenStudy (anonymous):

i'm not sure what to do after cosx=-1/2

OpenStudy (mertsj):

Are you familiar with the unit circle?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

And you know that the cosine is like x?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So find the two angles on the unit circle which have x coordinate - 1/2

OpenStudy (anonymous):

11pi/6 and 7pi/6?

OpenStudy (mertsj):